# Non-isothermal diffuse interface model for phase transition and interface evolution

### v1

@@ Abstract
In this paper, we derive a thermodynamically consistent non-isothermal diffuse interface model for phase transition and interface evolution involving heat transfer. This model is constructed by integrating concepts from classical irreversible thermodynamics with the energetic variational approach. By making specific assumptions about the kinematics of the temperature, we derive a non-isothermal Allen-Cahn equation. Through both asymptotic analysis and numerical simulations, we demonstrate that in the sharp interface limit, the non-isothermal Allen-Cahn equation converges to a two-phase Stefan-like problem, under a certain scaling of the melting/freezing energy. In this regime, the motion of the liquid-solid interface and the temperature interface coincide and are governed by the mean curvature, at least for a short time. The result provides a justification for the classical Stefan problem within a certain physical regime.
@@ 1 Introduction
Understanding the evolution of the interface between two phases undergoing a phase transition is an important problem in both physics and mathematics [1]. One classical approach to studying this behavior is through Stefan problems [45, 47, 49]. Generally speaking, the formulation of a Stefan problem involves solving a heat equation within each respective phase, while adhering to prescribed initial and boundary conditions. At the interface, where the two phases meet, the temperature is fixed at the critical temperature for phase transition - a key feature in classical Stefan problems. To complete the mathematical framework, an additional equation known as the Stefan condition is imposed. This condition, rooted in energy conservation principles, serves to precisely determine the position of the advancing interface. Stefan problems have been one of the most well-studied classes of free boundary problems dating back over a hundred years with the pioneering work by Lame [32] and Stefan [47] and more recent results by [45, 44, 5, 38, 42]. Key results such as the existence of solutions , the continuity of the moving boundary and the regularity of solutions can be found in a series of papers [8, 31, 10, 9, 3, 25]. However, it remains unclear whether such a simple model is truly valid in reality. For instance, classical Stefan problems typically consider only a single interface by fixing the temperature at the phase change temperature at the liquid-solid interface, which is not true in general. Whereas in reality, it is often observed that the liquid phase is below its freezing point or the solid phase is above it, which is known as supercooling and superheating [11, 46, 24, 4, 21]. More importantly, one could ask what is the thermodynamic reasoning behind Stefan problems? Are these model thermodynamically consistent, i.e., do they satisfy the first and second laws of thermodynamics?
@@
In this paper, we revisit this classical problem by developing a new thermodynamically consistent non-isothermal diffuse interface model for phase transition and interface evolution involving heat transfer. In contrast to the approaches in Stefan problems, which assume that there exists a sharp interface between the liquid and solid phase, diffuse interface models [2] assume that there exists an interfacial region between the two phases, which is closer to the physical reality [11]. One can introduce a smooth space-time-dependent phase function $\varphi({\bm{x}},t)$ to identify each phase, as well as the interface. Classical diffuse interface models often consider isothermal cases [2, 15]. However, to model phase transitions and heat transfer as considered in the Stefan problem [33], it is crucial to include the temperature as an additional variable, which is challenging because one requires the system to be both physically and mathematically consistent. Some advances in this direction have been made by Caginalp in a series of papers [11, 12, 14, 13], who was one of the first to study temperature dependent phase field models and who found the relation of these systems with free boundary problems, in particular the Stefan problem, as their sharp interface limit [13]. Around the same time, Penrose and Fife introduced their famous model, where both the phase parameter and the temperature can depend on space and time [43]. The connection between the Penrose-Fife model and Stefan-type problems has been explored in the literature [20, 19]. However, in both models, the temperature equation only contains linear terms in both the phase variable and temperature variable. In this work we take a new approach by combining ideas from the energetic variational approach with non-equilibrium thermodynamics. This approach works with the entropy of the system rather than the enthalpy as main thermodynamic quantity as most of the previous work, which allows us to include higher order nonlinearities of the system that could get lost in other cases. Furthermore, our approach allows for various assumptions regarding the kinematics of the temperature, corresponding to different heat transfer mechanisms in physical bodies. This flexibility enables applications across a wide range of physical models.
@@
The rest of paper is organized as follows: In Section 2, we derive, following a free energy approach, non-isothermal diffuse interface models, including the non-isothermal Allen-Cahn equation, under various assumptions on the kinematics of the temperature. In Section 3, we show that the formal asymptotic limit of the non-isothermal Allen-Cahn system, under certain scaling of the melting/freezing energy, leads to a two-phase Stefan-like problem with curvature effects. Hence, the Stefan problem can be seen as a reduced model of the non-isothermal Allen-Cahn equation. Finally, we perform some numerical studies to the non-isothermal Allen-Cahn model in Section 4.
@@ 2 Derivation of non-isothermal diffuse interface models
  • Formulate the free energy density $\psi(f,\theta)$ of the system, which includes both mechanical (resolved) and thermal (under-resolved) contributions. Here $f$ denotes the mechanical state variable, such as density and phase function, and $\theta$ represents the temperature. From this free energy, one can derive other thermodynamic quantities, such as entropy $s(f,\theta)$ and internal energy $e(f,\theta)$, according to the fundamental thermodynamic relations.
@@
  • Determine the equation of the temperature $\theta$ or entropy $s(f,\theta)$ by using the laws of thermodynamics and constitutive relations including the Clausius-Duhem relation and Fourier’s law.
@@ 2.1 Free energy formalism of diffuse interface models
Here, $\phi$ is the phase function, $\theta$ is the temperature, $\theta_{c}$ is the critical temperature for which the system undergoes the phase transition, $\epsilon$ is the small parameter related to the thickness of the interface., $W(\phi)$ is the interfacial potential energy. The parameter $c_{v}$ corresponds to the volumetric heat capacity, while $\chi(\epsilon)\lambda$ represents the parameters related to the latent heat of the system. We use $\chi(\epsilon)$ to denote the scaling of the latent heat with respect to the interfacial parameter $\epsilon$, which plays an important role in the following analysis. We refer to [40, 17] for details on the physical context of these parameters. It is worth mentioning that $\phi$ is a dimensionless variable that labels the two phases, and the free energy (2.1) is written in a non-dimensionalized form.
@@
The double-well potential can be seen as an approximation of the logarithmic potential. Both potentials have two global minima in the interior of the physically relevant domain $[-1,1]$, which is coherent with the physical interpretation of a diffuse-interface model in which only the values of the variable $\phi\in[-1,1]$ are meaningful. We will work with the double-well potential throughout this paper.
@@
Taking a closer look at the two additional terms in the free energy, when compared to the constant temperature case, we observe that the third term represents the melting/ freezing contribution to the energy. The idea is that for temperatures $\theta>\theta_{c}$ we assume that the joint potential of $W(\phi)-\chi(\epsilon)\frac{\lambda}{\theta_{c}}(\theta-\theta_{c})g(\phi)$ is no longer symmetric and that the local minimum for $\phi>0$ is smaller than the local minima for $\phi<0$. Thus, the global minimizer of the free energy is attained only for the state $\phi>0$. Similarly, we require a reverse statement to be true for $\theta<\theta_{c}$. Hence, we observe that the temperature moves the free energy in favor of one phase state. Two choices of $g(\phi)$ that satisfy the above are
@@ 2.2 Kinematics of the phase function and the temperature
Similar to the mechanical variable, we shall ask how the temperature $\theta$ is transported without any thermal-mechanical processes. We define $\frac{\tilde{D}}{\tilde{D}t}$ as the mechanical time derivative which describes the mechanical transport properties, known as the kinematics of temperature. We consider three possible situations, which lead to three distinct systems.
@@
    We assume that the temperature moves along with the material particles, i.e. $\theta$ is transported along the trajectory of the flow map with velocity $u$
@@
    We assume that the temperature is on a fixed background, i.e. $\theta$ is independent of the flow map with velocity $u$. This relation is equivalent to the temperature satisfying the equation
@@
    We assume that the temperature moves independent from the particles on the macroscopic scale, but is transported along the trajectory of another flow map, the macroscopic flow map $y(X,t)$, defined by a macroscopic background velocity $v$. For example, $v$ could represent the velocity in an incompressible Navier-Stokes equation. In this case
@@ 2.3 EnVarA and force balance
We start with the least action principle to derive the first contributions to the force balance equation. The least action principles states that the path taken by a Hamiltonian system between two states is the one for which the action functional $A$ is stationary (usually a minimum). This principle can be used to derive the the inertial and conservative forces in a dissipative system. More precisely, for a given initial volume $\Omega_{0}\subseteq\Omega$ (the reference configuration), its evolution at a specific time $t\in(0,T)$ is represented by $\Omega_{t}^{x}=\{z\in\Omega:z=x(X,t),~{}X\in\Omega_{0}\}$, where $x(X,t)$ is the flow map introduced in (2.6), and $\Omega_{t}^{x}$ is referred to as the current configuration. We can define the action $A$ as the integral of the kinetic energy minus the free energy over time $[0,T]$ as follows
$A(x(\cdot))=\int_{0}^{T}E^{kin}-\int_{\Omega_{t}^{x}}\psi(\phi,\nabla\phi,\theta)\,\textnormal{d}x\textnormal{d}t\ ,$
where $E^{kin}$ is the kinetic energy. One shall consider the action $A$ as a functional of the flow map $x(X,t)$ in Lagrangian coordinates, as will be done later in (2.8). The inertial and conservative forces, $f_{inertial}$ and $f_{cons}$, can be obtained by the variation of $A$ with respect to the flow map $x(X,t)$
@@
    $f_{inertial}=0,\quad f_{cons}=(\mu\nabla\phi+s\nabla\theta).$
@@
    $f_{inertial}^{v}=-(\partial_{t}v+v\cdot\nabla v)\quad f_{cons}^{v}=\nabla p$
@@
In this proof, we focus on case $(i)$ and only present the main differences for cases $(ii)$ and $(iii)$.
@@
Now, let assumption (A3) for the temperature hold. Since there are two flow maps $x(X,t)$ and $y(X,t)$ associated with the velocity $u$ and $v$ respectively, we have to apply the least action principle to both flow maps. For the flow map $x(X,t)$ moving with velocity $u$, we obtain the same contribution to the conservative forces as in case (A2) as $\theta$ is not transported by $u$. For the flow map $y(X,t)$ moving with velocity $v$, we obtain the following
$\delta A(y(\cdot))$ $=\int_{t\in[0,T]}\int_{\Omega_{t}^{y}}\{-(\partial_{t}v+v\cdot\nabla v)-\nabla p\}\cdot\delta y\,\textnormal{d}y\textnormal{d}t,$
where $p$ is the Lagrange multiplier for the incompressible condition, $\det F=1$.
@@
    $\partial_{t}v+(v\cdot\nabla)v+\nabla p$ $=-\mu\nabla\phi+\nu\Delta v$
    $\nabla\cdot v$ $=0.$
@@ 2.4 The temperature equation
What remains now is to derive the equation of the temperature $\theta$ or entropy $s$. The key ingredients are the first law of thermodynamics and the second law of thermodynamics, expressed by the Clausius-Duhem inequality.
The Clausius-Duhem inequality expresses the second law of thermodynamics, the increase of the total entropy, through the relation
@@
The Fourier law as an assumption on the heat flux is rather classical. Other choices such as a rescaled version where $q=k\frac{\nabla\theta}{\theta}$ can also be found in the literature, see for example [41]. Another choice is to assume a time dependent relation such as $\partial_{t}q+q=k\nabla\theta$, which is known as Cattaneo’s law [39].
To determine the form of $\Delta^{*}$, one need to use first law of thermodynamics, which states that the total energy for an isolated system, where no work is done on or by the system and no heat is exchanged with the environment, remains constant, i.e., $\frac{\textnormal{d}}{\textnormal{d}t}E^{tot}(t)=0$. This condition is typically specified by appropriate boundary conditions. To define the total energy we introduce the internal energy density $e=e(\phi,\theta)$ given by a Legendre transform of the free energy with respect to the temperature
@@
Using the boundary conditions for the heat flux q, we have
@@
as long as identity (2.16) is satisfied. This concludes the proof Theorem 2.7, part $(i)$. The argument to prove $(ii)$ under assumption (A2) is analogous.
@@
$=\int_{\Omega}((u-v)\cdot\nabla\phi\nabla\phi\eta(\phi,\theta)+\lambda\Delta v-\nabla p)\cdot v\,\textnormal{d}x$
@@
Combining the two integral identities we obtain
@@
$+\int_{\Omega}-v\cdot\nabla p+sv\cdot\nabla\theta+\nabla\cdot q\,\textnormal{d}x=0,$
where the last integrals equates to zero by using the incompressibility of $v$ and the boundary conditions for the system. Thus, the identity holds as long as identity (2.18) is satisfied. This concludes the proof of the theorem.

### v2

@@ Abstract
In this paper, we derive a thermodynamically consistent non-isothermal diffuse interface model for phase transition and interface evolution involving heat transfer. This model is constructed by integrating concepts from classical irreversible thermodynamics with the energetic variational approach. By making specific assumptions about the kinematics of the temperature, we derive a non-isothermal Allen-Cahn equation. Through both asymptotic analysis and numerical simulations, we demonstrate that in the sharp interface limit, the non-isothermal Allen-Cahn equation converges to a two-phase nonlinear Stefan type problem, under a certain scaling of the melting/freezing energy. In this regime, the motion of the liquid-solid interface and the temperature interface coincide and are governed by the mean curvature, at least for a short time. The result provides a justification for the classical Stefan problem within a certain physical regime.
@@ 1 Introduction
Understanding the evolution of the interface between two phases undergoing a phase transition is an important problem in both physics and mathematics [1]. One classical approach to studying this behavior is through Stefan problems [45, 47, 49]. Generally speaking, the formulation of a Stefan problem involves solving a heat equation within each respective phase, while adhering to prescribed initial and boundary conditions. At the interface, where the two phases meet, the temperature is fixed at the critical temperature for phase transition, which is a key feature of classical Stefan problems. To complete the mathematical framework, an additional equation known as the Stefan condition is imposed. This condition, rooted in energy conservation principles, serves to precisely determine the position of the advancing interface. Stefan problems have been one of the most well-studied classes of free boundary problems dating back over a hundred years with the pioneering work by Lame [32] and Stefan [47] and more recent results by [45, 44, 5, 38, 42]. Key results such as the existence of solutions , the continuity of the moving boundary and the regularity of solutions can be found in a series of papers [8, 31, 10, 9, 3, 25]. However, it remains unclear whether such a simple model is truly valid in reality. For instance, classical Stefan problems typically consider only a single interface by fixing the temperature at the phase change temperature at the liquid-solid interface, which is not true in general. Whereas in reality, it is often observed that the liquid phase is below its freezing point or the solid phase is above it, which is known as supercooling and superheating [11, 46, 24, 4, 21]. More importantly, one could ask what is the thermodynamic reasoning behind Stefan problems? Are these models thermodynamically consistent, i.e., do they satisfy the first and second laws of thermodynamics?
@@
In this paper, we revisit this classical problem by developing a new thermodynamically consistent non-isothermal diffuse interface model for phase transition and interface evolution involving heat transfer. In contrast to the approaches in Stefan problems, which assume that there exists a sharp interface between the liquid and solid phase, diffuse interface models [2] assume that there exists an interfacial region between the two phases, which is closer to the physical reality [11]. One can introduce a smooth space-time-dependent phase function $\phi({\bm{x}},t)$ to identify each phase, as well as the interface. Classical diffuse interface models often consider isothermal cases [2, 15]. However, to model phase transitions and heat transfer as considered in the Stefan problem [33], it is crucial to include the temperature as an additional variable, which is challenging because one requires the system to be both physically and mathematically consistent. Some advances in this direction have been made by Caginalp in a series of papers [11, 12, 14, 13], who was one of the first to study temperature dependent phase field models and who found the relation of these systems with free boundary problems, in particular the Stefan problem, as their sharp interface limit [13]. Around the same time, Penrose and Fife introduced their famous model, where both the phase parameter and the temperature can depend on space and time [43]. The connection between the Penrose-Fife model and Stefan-type problems has been explored in the literature [20, 19]. However, in both models, the temperature equation only contains linear terms in both the phase variable and temperature variable. In this work we take a new approach by combining ideas from the energetic variational approach with non-equilibrium thermodynamics. This approach works with the entropy of the system rather than the enthalpy as main thermodynamic quantity as most of the previous work, which allows us to include higher order nonlinearities of the system that could get lost in other cases. Furthermore, our approach allows for various assumptions regarding the kinematics of the temperature, corresponding to different heat transfer mechanisms in physical bodies. This flexibility enables applications across a wide range of physical models.
@@
The rest of paper is organized as follows: In Section 2, we derive, following a free energy approach, non-isothermal diffuse interface models, including the non-isothermal Allen-Cahn equation, under various assumptions on the kinematics of the temperature. In Section 3, we show that the formal asymptotic limit of the non-isothermal Allen-Cahn system, under certain scaling of the melting/freezing energy, leads to a nonlinear two-phase Stefan type problem. Hence, the Stefan problem can be seen as a reduced model of the non-isothermal Allen-Cahn equation. Finally, we perform some numerical studies to the non-isothermal Allen-Cahn model in Section 4.
@@ 2 Derivation of non-isothermal diffuse interface models
  • Formulate the free energy density $\psi(f,\theta)$ of the system, which includes both mechanical (resolved) and thermal (unresolved) contributions. Here $f$ denotes the mechanical state variable, such as density and phase function, and $\theta$ represents the temperature. From this free energy, one can derive other thermodynamic quantities, such as entropy $s(f,\theta)$ and internal energy $e(f,\theta)$, according to the fundamental thermodynamic relations.
@@
  • Determine the equation of the temperature $\theta$ or entropy $s(f,\theta)$ by using the laws of thermodynamics and constitutive relations, including the Clausius-Duhem relation and Fourier’s law.
@@ 2.1 Free energy formalism of diffuse interface models
Here, $\phi$ is the phase function, $\theta$ is the temperature, $\theta_{c}$ is the critical temperature at which the system undergoes the phase transition, $\epsilon$ is the small parameter related to the thickness of the interface, and $W(\phi)$ is the interfacial potential energy. The parameter $c_{v}$ corresponds to the volumetric heat capacity, while $\chi(\epsilon)\lambda$ represents the parameters related to the latent heat of the system. We use $\chi(\epsilon)$ to denote the scaling of the latent heat with respect to the interfacial parameter $\epsilon$, which plays an important role in the following analysis. We refer to [40, 17] for details on the physical context of these parameters. It is worth mentioning that $\phi$ is a dimensionless variable that labels the two phases, and the free energy (2.1) is written in a non-dimensionalized form.
@@
The double-well potential can be seen as an approximation of the logarithmic potential. Both potentials have two global minima in the interior of the physically relevant domain [1,1][-1,1][ - 1 , 1 ], which is coherent with the physical interpretation of a diffuse interface model in which only the values of the variable $\phi\in[-1,1]$ are meaningful. We will work with the double-well potential throughout this paper.
@@
Taking a closer look at the two additional terms in the free energy, when compared to the isothermal case, we observe that the third term represents the melting/freezing contribution to the free energy. The idea is that for temperatures $\theta>\theta_{c}$ we assume that the joint potential of $W(\phi)-\chi(\epsilon)\frac{\lambda}{\theta_{c}}(\theta-\theta_{c})g(\phi)$ is no longer symmetric and that the local minimum for $\phi>0$ is smaller than the local minima for $\phi<0$. Thus, the global minimizer of the free energy is attained only for the state $\phi>0$. Similarly, we require a reverse statement to be true for $\theta<\theta_{c}$. Then, the temperature moves the free energy in favor of one phase state. Two choices of $g(\phi)$ that satisfy the above are
@@ 2.2 Kinematics of the phase function and the temperature
Similar to the mechanical variable, we shall ask how the temperature $\theta$ is transported without any thermal-mechanical processes. We define $\frac{\tilde{D}}{\tilde{D}t}$ as the mechanical time derivative, which describes the mechanical transport properties, known as the kinematics of temperature. We consider three possible situations, which lead to three distinct systems:
@@
    We assume that the temperature moves along with the material particles, i.e. $\theta$ is transported along the trajectory of the flow map with velocity $u$, given by
@@
    We assume that the temperature is on a fixed background, i.e. $\theta$ is independent of the flow map with velocity $u$. This relation is equivalent to the temperature satisfying the equation:
@@
    We assume that the temperature moves independent from the particles on the macroscopic scale, but is transported along the trajectory of another flow map, the macroscopic flow map $y(X,t)$, defined by a macroscopic background velocity $v$. For example, $v$ could represent the velocity in an incompressible Navier-Stokes equation. In this case,
@@ 2.3 EnVarA and force balance
[…]
$A(x(\cdot))=\int_{0}^{T}\left(E^{kin}-\int_{\Omega_{t}^{x}}\psi(\phi,\nabla\phi,\theta)\,\textnormal{d}x\right)\textnormal{d}t\ ,$
where $E^{kin}$ is the kinetic energy. One shall consider the action $A$ as a functional of the flow map $x(X,t)$ in Lagrangian coordinates, as will be done later in (2.8). The inertial and conservative forces, $f_{inertial}$ and $f_{cons}$, can be obtained by the variation of $A$ with respect to the flow map $x(X,t)$, i.e.,
@@
    $f_{inertial}=0,\quad f_{cons}=\mu\nabla\phi+s\nabla\theta.$
@@
    $f_{inertial}^{v}=-(\partial_{t}v+v\cdot\nabla v)\quad f_{cons}^{v}=\nabla p+s\nabla\theta$
@@
[…]
@@
Now, let assumption (A3) for the temperature hold. Since there are two flow maps $x(X,t)$ and $y(X,t)$ associated with the velocity $u$ and $v$ respectively, we have to apply the least action principle to both flow maps. For the flow map $x(X,t)$ moving with velocity $u$, we obtain the same contribution to the conservative forces as in case (A2) as $\theta$ is not transported by $u$. However, as the temperature is transported by $v$ we obtain for the flow map $y(X,t)$ moving with velocity $v$ that
$\delta A(y(\cdot))$ $=\int_{t\in[0,T]}\int_{\Omega_{t}^{y}}\{-(\partial_{t}v+v\cdot\nabla v)-\nabla p+s\nabla\theta\}\cdot\delta y\,\textnormal{d}y\textnormal{d}t,$
where $p$ is the Lagrange multiplier for the incompressible condition, $\det F=1$. The last term follows from the fact that the free energy under the flow $y$ only depends on the temperature $\theta$.
@@
    $\partial_{t}v+(v\cdot\nabla)v+\nabla p$ $=-\mu\nabla\phi+\nu\Delta v-s\nabla\theta$
    $\nabla\cdot v$ =0=0.= 0 .
@@ 2.4 The temperature equation
What remains now is to derive the equation of the temperature $\theta$ or entropy $s$. The derivation relies on the first and second laws of thermodynamics, with the latter expressed through the Clausius-Duhem inequality.
The Clausius-Duhem inequality provides a mathematical expression of the second law of thermodynamics, the increase of the total entropy. It can be written in the following equivalent forms:
@@
The Fourier law is a classical assumption on the heat flux. Alternative formulations have also been proposed in the literature. For instance, a rescaled version, $q=k\frac{\nabla\theta}{\theta}$, can be found in the literature, see for example [41]. Another choice is to assume a time dependent relation $\partial_{t}q+q=k\nabla\theta$, which is known as Cattaneo’s law [39].
To determine the form of $\Delta^{*}$, one need to use first law of thermodynamics, which states that the total energy for an isolated system, where no work is done on or by the system and no heat is exchanged with the environment, remains constant, i.e., $\frac{\textnormal{d}}{\textnormal{d}t}E^{tot}(t)=0$. To enforce this condition, we assume that the heat flux satisfies a non-flux boundary condition, i.e.,
(2.14) $q\cdot n=0.$
To define the total energy we introduce the internal energy density $e=e(\phi,\theta)$ given by a Legendre transform of the free energy with respect to the temperature
@@
Using the boundary conditions for the heat flux $q$ (2.14), we have
@@
[…]
@@
$=\int_{\Omega}((u-v)\cdot\nabla\phi\nabla\phi\eta(\phi,\theta)+\nu\Delta v-\nabla p-s\nabla\theta)\cdot v\,\textnormal{d}x.$
@@
Combining the two integral identities and using integration by parts, we obtain
@@
$+\int_{\Omega}-v\cdot\nabla p+\nabla\cdot q\,\textnormal{d}x=0,$
where the last integral vanishes due to the incompressibility of $v$, the non-flux boundary condition for the heat flux (2.14), and the boundary condition $v\cdot n=0$. Thus, the identity holds as long as identity (2.19) is satisfied. This concludes the proof of the theorem.